Selling price per unit $               120
Current annual sales $        240,000
Annual fixed costs
Factory rent $          43,400
Depreciation expense - equipment $          12,000
Utilities $          22,000
Insurance $            8,400
$     85,800
Variable costs
Direct materials $                 36 per unit
Direct labour $                 48 per unit
Income tax rate 20%

If Bridgeport would like to earn a profit after tax of $13,000, what should the sales be?

At this sales level, what is the degree of operating leverage?
What is the margin of safety in unit?
Answer #1
$
Selling price per unit 120
Less: Variable cost
Direct material 36
Direct labor 48
Contribution margin per unit 36
a. Profit after tax = $ 13000
Rate of tax is 20%
Profit before tax = $ 13000 * 100 / 80 = $ 16250
Proft before tax + fixed cost = Total Contribution margin
$ 16250 + $ 85800 = $ 102050
No. of units to be sold = Total contribution margin / contribution margin per unit
No. of units to be sold = $ 102050 / $ 36 = 2835 units
Sales (in $ ) = 2835 units * $ 120 = $ 340200
Operating leverage = Quantity x (sale price per unit - variable cost per unit)
Quantity x (sale price per unit - variable cost per unit) - fixed operating cost
Operating leverage = 2835 X (120 - 36) = 238140 = 1.56
2835 X (120 - 36) - 85800 152340
Breakeven sale (in units) = Fixed cost / contribution margin per unit = $ 85800 / $ 36 = 2383 units
Margin of safety (in units) = Projected sale - breakeven sale
Margin of safety (in units) = 2835 - 2383 = 452 units